# Primordial Black Holes from Kinetic Preheating

### v1

@@ Abstract
We demonstrate that violent kinetic preheating following inflation can lead to the formation of black holes in the early Universe. In $\alpha$-attractor models with derivative inflaton couplings, nonlinear amplification of field fluctuations drives large spacetime curvature and gravitational collapse shortly after inflation ends. Using fully general-relativistic lattice simulations, we find that these dynamics produce black holes with masses of order tens of grams at sub-horizon scales, without requiring large primordial curvature perturbations. Although such micro-black holes evaporate rapidly via Hawking radiation, their formation modifies the post-inflationary equation of state and their evaporation can successfully reheat the Universe before Big Bang nucleosynthesis. These results identify kinetic preheating as a new, efficient channel for black-hole production and establish a direct connection between inflationary symmetries and strong-gravity phenomena at reheating.
@@ I Introduction
The inflationary epoch not only explains the observed large-scale homogeneity of the Universe Guth (1981); Albrecht and Steinhardt (1982); Linde (1983), but also provides a natural origin for primordial fluctuations that seed cosmic structure Guth and Pi (1982); Hawking (1982); Bardeen et al. (1983). Yet, the transition from inflation to the hot Big Bang, reheating, remains one of the least understood phases in the early Universe. In many well-motivated models, the inflaton’s couplings to matter fields can drive non-perturbative and highly non-linear energy transfer, a process known as preheating Traschen and Brandenberger (1990); Shtanov et al. (1995); Kofman et al. (1994, 1997); Greene et al. (1997). Preheating generically leads to the production of a high-frequency (MHz-GHz) stochastic gravitational wave background due to the generation of large gradients in the energy density Khlebnikov and Tkachev (1997); Easther and Lim (2006); Easther et al. (2007); Garcia-Bellido and Figueroa (2007); Garcia-Bellido et al. (2008); Dufaux et al. (2007); Easther et al. (2008); Bethke et al. (2014); Dufaux et al. (2010); Garcia-Bellido and Figueroa (2007); Figueroa et al. (2016); Figueroa and Torrenti (2017). When the transfer proceeds through derivative, or kinetic couplings rather than potential-like interactions, the resulting dynamics can be especially violent Adshead et al. (2015); Cuissa and Figueroa (2019); Adshead et al. (2024a), sourcing large gradients Adshead et al. (2016), strong gravitational waves Adshead et al. (2018, 2020a, 2020b, 2024b, 2024c), and as we demonstrate in this letter, gravitational collapse to black holes.
@@
Unlike primordial black holes (PBHs) Hawking (1971); Carr and Hawking (1974); Carr (1975); Green and Liddle (1997); Musco et al. (2005); Musco and Miller (2013); Harada et al. (2013) formed from large curvature perturbations during inflation Garcia-Bellido and Ruiz Morales (2017); Byrnes et al. (2019); Sasaki et al. (2018), the black holes produced here originate from the intrinsically nonlinear, post-inflationary dynamics of preheating. Their formation reflects the strong coupling between field gradients and spacetime curvature in the fully relativistic regime, which we capture using GABERel Giblin and Tishue (2019); Adshead et al. (2024b) — an extension of GABE Child and Giblin (2012) that evolves the metric and matter fields self-consistently in the Baumgarte-Shapiro-Shibata-Nakamura (BSSN) scheme of numerical relativity Baumgarte and Shapiro (1998); Shibata and Nakamura (1995). The resulting gravitational collapse occurs at sub-horizon scales and within a few oscillations of the inflaton field, highlighting the extreme efficiency of energy localization during kinetic preheating.
The black holes produced in this scenario are far too light to survive to the present day, evaporating via Hawking radiation shortly after formation. Nevertheless, their transient existence can have significant cosmological implications. Their evaporation is sufficient to reheat the Universe without any further couplings between the inflationary sector and the standard model sector. Further, a PBH dominated phase may lead to nonthermal particle production, modify the post-inflationary equation of state, or imprint characteristic features in the stochastic gravitational-wave spectrum. More broadly, these results establish kinetic preheating as a qualitatively new channel for black-hole formation in the early Universe — one that connects the symmetry structure of inflationary models directly to gravitational collapse and quantum gravity–scale phenomena.
@@ IV Reheating from Primordial Black-Hole Evaporation
Subsequently a matter-dominated phase follows until the black holes evaporate via Hawking radiation Hawking (1975); Page (1976); Anantua et al. (2009) at
@@
so that $M_{\rm BH}\simeq 30\,{\rm g}$ yields $T_{\rm reh}\sim 10^{8}\,{\rm GeV}$. Because $t_{\rm evap}\ll 1\,{\rm s}$, this reheating occurs well before Big-Bang nucleosynthesis, leaving standard light-element abundances unaffected. Hawking emission distributes roughly $f_{g}\sim 10^{-2}$ Page (1976) of the PBH energy into gravitons, giving a negligible $\Delta N_{\rm eff}^{\rm(GW)}\lesssim 10^{-3}$, while any additional decoupled light species that receives a fraction of the energy, [^1] Here, $g_{X}$ and $g_{\star}^{\rm vis}$ are the effective numbers of relativistic degrees of freedom in the additional species, $X$, and the visible standard model sector, respoectively. $f_{X}\sim g_{X}/(g_{X}+g_{\star}^{\rm vis})$, contributes $\Delta N_{\rm eff}\simeq 0.1(f_{X}/0.03)$. In particular, three right-handed (sterile) neutrinos or other light fermions added to the Standard Model contribute only $\Delta N_{\rm eff}\approx 0.1$, comfortably below present CMB bounds — although within the range probed by future experiments Abazajian et al. (2016, 2019, 2022).
Thus, evaporation of the micro-black holes formed during kinetic preheating provides a natural and efficient reheating mechanism: it restores a hot radiation bath at $T_{\rm reh}\sim 10^{10-12}{\rm GeV}$, precedes BBN by many orders of magnitude, and leaves only a minuscule residual contribution to relativistic energy density today. Kinetic preheating after $\alpha$-attractor inflation therefore provides a realization of the scenarios discussed in Riajul Haque et al. (2023).

### v2

@@ Abstract
We demonstrate that violent kinetic preheating following inflation can lead to the formation of black holes in the early Universe. In $\alpha$-attractor models with derivative inflaton couplings, nonlinear amplification of field fluctuations drives large spacetime curvature and gravitational collapse shortly after inflation ends. Using fully general-relativistic lattice simulations, we find that these dynamics produce black holes with masses of order tens of grams at sub-horizon scales, without requiring large primordial curvature perturbations. Although such micro-black holes evaporate rapidly via Hawking radiation, their formation modifies the post-inflationary equation of state and their evaporation can successfully reheat the Universe before Big Bang nucleosynthesis. These results identify kinetic preheating as a new, efficient channel for black-hole production.
@@ I Introduction
The inflationary epoch not only explains the observed large-scale homogeneity of the Universe Guth (1981); Albrecht and Steinhardt (1982); Linde (1983), but also provides a natural origin for primordial fluctuations that seed cosmic structure Guth and Pi (1982); Hawking (1982); Bardeen et al. (1983). Yet, the transition from inflation to the hot Big Bang, reheating, remains one of the least understood phases in the early Universe. In many well-motivated models, the inflaton’s couplings to matter fields can drive non-perturbative and highly non-linear energy transfer, a process known as preheating Traschen and Brandenberger (1990); Shtanov et al. (1995); Kofman et al. (1994, 1997); Greene et al. (1997). Preheating generically leads to the production of a high-frequency (MHz-GHz) stochastic gravitational wave background due to the generation of large gradients in the energy density Khlebnikov and Tkachev (1997); Easther and Lim (2006); Easther et al. (2007); Garcia-Bellido and Figueroa (2007); Garcia-Bellido et al. (2008); Dufaux et al. (2007); Easther et al. (2008); Bethke et al. (2014); Dufaux et al. (2010); Garcia-Bellido and Figueroa (2007); Figueroa et al. (2016); Figueroa and Torrenti (2017). When the transfer proceeds through derivative, or kinetic couplings rather than potential-like interactions, the resulting dynamics can be especially violent Adshead et al. (2015); Cuissa and Figueroa (2019); Adshead et al. (2024a), sourcing large gradients Adshead et al. (2016), strong gravitational waves Adshead et al. (2018, 2020a, 2020b); Weiner:2020sxn; Adshead et al. (2024b, c), and as we demonstrate in this letter, gravitational collapse to black holes.
@@
Unlike primordial black holes (PBHs) Hawking (1971); Carr and Hawking (1974); Carr (1975); Green and Liddle (1997); Musco et al. (2005); Khlopov (2010); Musco and Miller (2013); Harada et al. (2013) formed from large curvature perturbations during inflation Garcia-Bellido and Ruiz Morales (2017); Byrnes et al. (2019); Sasaki et al. (2018); Bhattacharya et al. (2020); Martin et al. (2020), the black holes produced here originate from the intrinsically nonlinear, post-inflationary dynamics of preheating. Their formation reflects the strong coupling between field gradients and spacetime curvature in the fully relativistic regime, which we capture using GABERel Giblin and Tishue (2019); Adshead et al. (2024b) — an extension of GABE Child and Giblin (2012) that evolves the metric and matter fields self-consistently in the Baumgarte-Shapiro-Shibata-Nakamura (BSSN) scheme of numerical relativity Baumgarte and Shapiro (1998); Shibata and Nakamura (1995). The resulting gravitational collapse occurs at sub-horizon scales and within a few oscillations of the inflaton field, highlighting the extreme efficiency of energy localization during kinetic preheating.
The black holes produced in this scenario are far too light to survive to the present day, evaporating via Hawking radiation shortly after formation. Nevertheless, their transient existence can have significant cosmological implications. Their evaporation is sufficient to reheat the Universe without any further couplings between the inflationary sector and the standard model sector. Further, a PBH dominated phase may lead to nonthermal particle production, modify the post-inflationary equation of state, or imprint characteristic features in the stochastic gravitational-wave spectrum. More broadly, these results establish kinetic preheating as a qualitatively new channel for black-hole formation in the early Universe.
@@ IV Reheating from Primordial Black-Hole Evaporation
Subsequently a matter-dominated phase follows until the black holes evaporate via Hawking radiation Hawking (1975); Page (1976); Anantua et al. (2009); Zagorac:2019ekv at
@@
so that $M_{\rm BH}\simeq 30\,{\rm g}$ yields $T_{\rm reh}\sim 10^{8}\,{\rm GeV}$. Because $t_{\rm evap}\ll 1\,{\rm s}$, this reheating occurs well before Big-Bang nucleosynthesis, leaving standard light-element abundances unaffected. Hawking emission distributes roughly $f_{g}\sim 10^{-2}$ Page (1976) of the PBH energy into gravitons, giving a negligible $\Delta N_{\rm eff}^{\rm(GW)}\lesssim 10^{-3}$. However, any additional decoupled light species attains approximately the same temperature as the visible Standard Model, and therefore contributes $\Delta N_{\rm eff}\simeq 0.027(g^{b}_{X}+7g^{f}_{X}/8)$, where $g^{b,f}_{X}$ counts the internal degrees of freedom for bosons or (Weyl) fermions, respectively. In particular, an additional three right-handed (sterile) neutrinos added to the Standard Model contribute $\Delta N_{\rm eff}\approx 0.14$. While this is within current bounds, it is well within the range targeted by upcoming experiments Abazajian et al. (2016, 2019, 2022).
Thus, evaporation of the micro-black holes formed during kinetic preheating provides a natural and efficient reheating mechanism: it restores a hot radiation bath at $T_{\rm reh}\sim 10^{8-10}{\rm GeV}$, precedes BBN by many orders of magnitude, and leaves only a minuscule residual contribution to relativistic energy density today. Kinetic preheating after $\alpha$-attractor inflation therefore provides a realization of the scenarios discussed in Riajul Haque et al. (2023).